# Linear Singer-Hopf conjecture

### v1

@@ Abstract.
If $X$ is a closed $2n$-dimensional aspherical manifold, i.e., the universal cover of $X$ is contractible, then the Singer-Hopf conjecture predicts that $(-1)^{n}\chi(X)\geq 0$. We prove this conjecture when $X$ is a complex projective manifold whose fundamental group admits an almost faithful linear representation over any field. In fact, we prove a much stronger statement that if $X$ is a complex projective manifold with large fundamental group and $\pi_{1}(X)$ admits an almost faithful linear representation, then $\chi(X,\mathcal{P})\geq 0$ for any perverse sheaf $\mathcal{P}$ on $X$.
To prove the main result, we introduce a vanishing cycle functor of multivalued one-forms. Then using techniques from non-abelian Hodge theories in both archimedean and non-archimedean settings, we deduce the desired positivity from the geometry of pure and mixed period maps.
@@ 1. Introduction
In the 1930s, Hopf conjectured that a closed $2n$-dimensional compact Riemannian manifold $X$ with non-positive sectional curvature satisfies $(-1)^{n}\chi(X)\geq 0$ (see [Yau82, Problem 10]). Singer proposed to study Hopf’s conjecture using $L^{2}$-cohomology groups of the universal cover of $X$, which naturally lead to the following more general conjecture.
Conjecture 1.1 (Singer-Hopf).
Let $X$ be a closed $2n$-dimensional manifold. If $X$ is aspherical, i.e., the universal cover of $X$ is contractible, then
@@
The above conjecture was also considered by Thurston for 4-manifolds ([Kir97, Problem 4.10]). When $n=1$, 1.1 follows from the easy fact that a Riemann surface is aspherical if and only if its genus is at least 1. If we know that $X$ admits a Riemannian metric with non-positive sectional curvature, then the conjecture follows from the Gauss-Bonnet theorem when $n=2$ (see [Che55]). Beyond this case, the conjecture is widely open.
If we assume $X$ is a complex projective manifold, then some progresses have been made in [Gro91, JZ00, CX01, DCL24, LMW21, AW21, LIP24], to quote only a few. In [LMW21], it was noticed that for projective manifolds, a natural generalization of being aspherical is having large fundamental group. Here we recall that a projective manifold $X$ is said to have large fundamental group if for any irreducible subvariety $Z\subset X$, the image of the composition $\pi_{1}(Z_{\operatorname{norm}})\to\pi_{1}(Z)\to\pi_{1}(X)$ is infinite, where $Z_{\operatorname{norm}}$ denotes the normalization of $Z$. Moreover, the assertion that $(-1)^{n}\chi(X)\geq 0$ was strengthened to
@@
These ideas were further pursued in [AW21]. Even though not explicitly stated, the following conjecture was expected to hold.
@@
In [AW21], the above conjecture is proved when there exists a semisimple, almost faithful, cohomologically rigid representation $\pi_{1}(X)\to\mathrm{GL}(r,\mathbb{C})$. Here we recall that a linear representation is called almost faithful if the kernel of the representation is finite, and it is called cohomologically rigid, if it has no nontrivial first order deformation.
In this paper, we prove 1.2 within a much broader framework: we merely require the existence of an almost faithful linear representation $\pi_{1}(X)\to\mathrm{GL}(r,K)$ over a field $K$ of arbitrary characteristic, without assuming semisimplity or cohomological ridigity of the representation.
@@
Let $X$ be a projective manifold. If there is a large representation $\rho:\pi_{1}(X)\to\mathrm{GL}(r,K)$ where $K$ is any field, then $\chi(X,{\mathcal{P}})\geq 0$ for any perverse sheaves ${\mathcal{P}}$ on $X$. In particular, $(-1)^{\dim X}\chi(X)\geq 0$.
Notice that if $X$ has large fundamental group, and if $\rho$ is almost faithful, then $\rho$ is large. Therefore, the following corollary is immediate.
@@
Let $X$ be a projective manifold with large fundamental group (e.g. $X$ is aspherical). If there exists an almost faithful representation $\rho:\pi_{1}(X)\to\mathrm{GL}(r,K)$, then for any perverse sheaf ${\mathcal{P}}$ on $X$, $\chi(X,{\mathcal{P}})\geq 0$. In particular, Singer-Hopf conjecture holds for $X$.
A slightly stronger formulation of Theorem 1.3 is the following.
@@ 1.1. Sketch of the proof
We will use techniques from the linear Shafarevich conjecture in [DYK23, DY24, Eys04, EKPR12] to establish Theorem 1.5. As in [AW21], Arapura and the second named author proved Theorem 1.5 assuming $\rho$ underlies a complex variation of Hodge structure ($\mathbb{C}$-VHS for short) with discrete monodromy. When $\rho$ does not satisfy such property, in [DYK23] Yamanoi and the first named author showed that one can produce non-trivial multi-valued one-forms on $X$. Our approach involves an iterative utilization of both multi-valued one-forms and the period maps of $\mathbb{C}$-VHS to deduce the inequality (2).
Let us briefly explain our strategy. We first observe that if there is a $d$-valued one-form whose image in $T^{*}X$ intersects $T^{*}_{Z}X$ at finitely many points, then the inequality (2) holds, since $d$ times the left-hand side is equal to the number of intersection points counting multiplicity.
In general, we introduce a vanishing cycle functor $\Phi_{\eta}$ of a $d$-valued one-form $\eta$ acting on the free abelian group generated by conormal varieties. We show that $d\,T^{*}_{Z}X\cdot T^{*}_{X}X=\Phi_{\eta}(T^{*}_{Z}X)\cdot T^{*}_{X}X$. Using $\Phi_{\eta}$, we can reduce (2) to the same inequality with $Z$ replaced by smaller subvarieties, as long as the restriction of $\eta$ to the smooth locus of $Z$ is non-trivial. Then the proof of Theorem 1.5 is divided into the following three cases.
Case (i): when $\mathrm{char}(K)>0$. In this case, we can prove that the above reduction process continues until $Z$ becomes points. When $Z$ is a point, the inequality (2) is obvious.
Case (ii): when $K=\mathbb{C}$ and $\rho$ is semisimple. In this case, when the above reduction process terminates, $Z_{\operatorname{norm}}$ underlies a $\mathbb{C}$-VHS with large monodromy representation and discrete monodromy group. As proved in [AW21], $Z_{\operatorname{norm}}$ then admits a finite morphism to the period domain of such $\mathbb{C}$-VHS. Then the inequality (2) can be deduced from the curvature property of the period domain.
Case (iii): when $K=\mathbb{C}$ and $\rho$ is not semisimple. In this case, when the reduction process terminates, $Z_{\operatorname{norm}}$ admits a complex variation of mixed Hodge structure ($\mathbb{R}$-VMHS) such that its mixed period map $Z_{\rm norm}^{\rm univ}\to\mathscr{M}$ has discrete fibers. In this case, two extra difficulties occur compared to Case (ii):
@@
  • The mixed period domain does not have the desired non-positive curvature in the pure case.
This is resolved by a technical result (cf. Proposition 5.3) which allows us to simultaneously explore the geometry of the mixed period map of the $\mathbb{R}$-VMHS and the period map of the $\mathbb{C}$-VHS corresponding to the semi-simplification of the $\mathbb{R}$-VMHS.
In [AW21], we obtain discreteness of the monodromy group from a deep theorem of Esnault-Groechenig ([EG18], see also [Esn23] for a survey). In this paper, the discreteness of monodromy groups are deduced from nonarchemidean Hodge theory, which is same as the approach in [BKT13].
In theory, we could bypass Case (ii) and directly prove the more general Case (iii). However, we decide to write the proof in separate cases to isolate the few technical arguments and to benefit the readers who would be satisfied understanding the proof up to Case (ii).
@@ 1.2. Relation with the Shafarevich conjecture and hyperbolicity
    when $X$ is not necessarily smooth and $\pi_{1}(X)$ admits a faithful reductive linear representation over $\mathbb{C}$, by the first named author, Yamanoi and Katzarkov in [DYK23];
@@
The proofs of the Shafarevich conjecture use nonabelian Hodge theories both in the archimedean setting by Simpson [Sim92] and non-archimedean setting by Gromov-Schoen [GS92].
In this paper, we do not use any result about holomorphic convexity or Steinness. Nevertheless, we use both archimedian and nonarchimedian nonabelian Hodge theories in the same way as in the proof of Shafarevich conjectures. These techiniques are robust tools in studying the Shafarevich conjecture, hyperbolicity of algebraic varieties (cf. e.g. [Yam10, CDY22, DY24]), and the Singer-Hopf conjecture in the linear case. Our main result Theorem 1.5 is certain positivity property of the cotangent bundle of $X$, which is known to be related to the hyperbolicity of $X$. For example, a projective manifold with ample cotangent bundle is Kobayashi hyperbolic.
Even though our result relies on $X$ to have a large fundamental group, we hope our positivity (2) can be directly related to hyperbolicity. For example, does (2) always hold for irreducible subvarieties $Z$ in a Kobayashi hyperbolic projective manifold $X$? In general, it is also interesting to find sufficient conditions on a representation of $\pi_{1}(X)$ such that (2) is always strict inequality. In [CDY22, DY24], it is proved that when the Zariski closure of a (generically) large representation $\varrho:\pi_{1}(X)\to\mathrm{GL}(r,K)$ is a semisimple algebraic group, then $X$ is “almost” hyperbolic. This condition is sharp as one has to exclude the case of abelian varieties. Hence we conjecture that the inequality (2) is strict if the Zariski closure of $\varrho({\rm Im}[\pi_{1}(Z_{\rm norm})\to\pi_{1}(X)])$ is a semisimple algebraic group for every positive-dimensional subvariety $Z$ of $X$. In this case, $X$ is Kobayashi hyperbolic by [Yam10, CDY22, DY24].
Recently, various generalizations of the Singer-Hopf conjecture in the algebra-geometric setting are formulated, e.g., in the singular setting as in [Max23], and in the coherent setting as in [AMW23]. Here, we also want to propose Singer-Hopf type conjectures for quasi-projective varieties. For example, we conjecture that if the universal cover of a quasi-projective manifold $X$ is a bounded symmetric space, then $\chi(X,{\mathcal{P}})\geq 0$ for any perverse sheaf on $X$ (with respect to an algebraic stratification).
We propose the following notion of “topological hyperbolicity” of algebraic varieties, aiming to capture analogous hyperbolic properties of Kähler hyperbolic manifolds.
@@ Acknowledgement
This paper builds upon previous collaborations of the second author with Yongqiang Liu, Laurentiu Maxim, and Donu Arapura, as well as the collaboration of the first author with Katsutoshi Yamanoi. We are grateful of the numerous enlightening conversations that have played a crucial role in the development of this work. We thank Jörg Schürmann for carefully reviewing an earlier draft of this paper and providing valuable feedback. We also thank Hélène Esnault for patiently answering our questions. Additionally, we have benefited greatly from discussions with Xiping Zhang, Jie Liu, and Qizheng Yin. The second author thanks Peking University and BICMR for the generous hospitality and support during the writing of this paper. The first author is partially supported by the French Agence Nationale de la Recherche (ANR) under reference ANR-21-CE40-0010 (KARMAPOLIS). The second author is partially supported by a Simons fellowship.
@@ 2. Preliminaries
We first review some results about constructible functions, constructible sheaves, characteristic cycles and the global index formula. The recent survey article [MS22] gives a more detailed summery of these subjects.
@@ 2.1. Constructible functions
is well-defined when the above stratification can be chosen to be finite and each stratum is homotopy equivalent to a finite CW-complex, e.g., when $X$ is compact or when $X$ is an algebraic variety and each stratum is a locally closed algebraic subvariety.
@@ 2.2. Constructible sheaves and characteristic cycles
where $K_{0}(D_{c}^{b}(X,K))$ is the Grothendieck group of $D_{c}^{b}(X,K)$. We will not review its precise definition (see e.g., [Dim04, Definition 4.3.19] and [KS90, Chapter IX]). Instead, we focus on the following two important properties.
@@
Note that even though $T^{*}X$ is not compact, the intersection number is well-defined because it can be defined as a zero cycle in the support of ${\mathcal{F}}$ in $X$ (see [Ful98, Chapter 6]).
@@ 2.4. Multivalued one-forms
Definition 2.7 (Mutivalued one-form).
Let $X$ be a complex manifold and let $E$ be a holomorphic vector bundle on $X$. A multivalued section $\eta$ on $X$ is a formal sum $\sum_{i=1}^{m}n_{i}\Gamma_{i}$, where each $n_{i}\in\mathbb{Z}_{>0}$ and each $\Gamma_{i}$ is an irreducible closed subvariety of $E$ such that the natural projection $\Gamma_{i}\to X$ is a finite surjective morphism. Such multisection $\eta$ is called $d$-valued for $d=\sum_{i=1}^{m}n_{i}\deg[\Gamma_{i}:X]$. We say that $\eta$ is reduced if all $\Gamma_{i}$ are distinct, and $\eta$ is irreducible if $m=1$ and $n_{1}=1$. We say that $\eta$ is trivial if $m=1$ and $\Gamma_{1}$ is the zero section of $E\to X$. Multivalued sections of $T^{*}X$ will be called mutivalued (holomorphic) one-forms.
Let $\eta$ be a mutivalued one-form on a complex manifold $X$. We sometimes denote by $\Gamma_{\eta}$ instead of $\eta$ to emphasize it is an analytic cycle in $T^{*}X$. By definition, there exists a largest Zariski open subset $X^{\circ}\subset X$ such that $\Gamma_{\eta}\cap T^{*}X^{\circ}\to X^{\circ}$ is étale. We call $X^{\circ}$ to be the unbranching locus of $\eta$ and the complement $X\setminus X^{\circ}$ the branching locus. In this case, for any $x\in X^{\circ}$, there exist an open neighborhood of $U$ of $x$ and holomorphic one-forms $\eta_{1},\ldots,\eta_{m}$ on $U$ such that $\eta|_{U}$ is equal to the union of $\eta_{1},\ldots,\eta_{m}$. Therefore, we say that $\eta$ is closed if for any point $x\in X^{\circ}$, each of the above $\eta_{i}$ is a closed one-form.
@@
Without loss of generality, we can assume that $\eta$ is irreducible. Since $X$ is projective and the natural map $\Gamma_{\eta}\to X$ is finite, $\Gamma_{\eta}$ is a projective variety. Let $Y\to\Gamma_{\eta}$ be a resolution of singularity, which is an isomorphism over the smooth locus of $\Gamma_{\eta}$. Let $\theta$ be the tautological holomorphic one form on $T^{*}X$, and let $\theta_{Y}$ be the pullback of $\theta$ via the composition $Y\to\Gamma_{\eta}\to T^{*}X$. Since $Y$ is a projective manifold, by Hodge theory $\theta_{Y}$ must be closed. Over the unbranching locus of $\eta$, the composition $Y\to\Gamma_{\eta}\to X$ is étale, and the pushforward of $\theta_{Y}$ to $X$ is equal to $\eta$. Thus, $\eta$ is closed. ∎
@@
We will call the above map $s_{\rho}$ the (Katzarkov-Eyssidieux) reduction map for $\rho$. The construction of the mutivalued holomorphic one-form $\eta_{\rho}$ associated with $\rho$ can be found in [CDY22, Step 2 in the proof of Theorem H]. The closedness of $\eta_{\rho}$ follows from Lemma 2.9. The equivalence of Item 1 and Item 3 is proved by Katzarkov [Kat97] and Eyssidieux [Eys04]. The equivalence of the first three items can be found in [CDY22, Theorem H]. Hence we only need to prove their equivalence to Item 4.

### v2

# Linear Chern-Hopf-Thurston conjecture

@@ Abstract.
If $X$ is a closed $2n$-dimensional aspherical manifold, i.e., the universal cover of $X$ is contractible, then the Chern-Hopf-Thurston conjecture predicts that $(-1)^{n}\chi(X)\geq 0$. We prove this conjecture when $X$ is a complex projective manifold whose fundamental group admits an almost faithful linear representation over any field. In fact, we prove a much stronger statement that if $X$ is a complex projective manifold with large fundamental group and $\pi_{1}(X)$ admits an almost faithful linear representation, then $\chi(X,\mathcal{P})\geq 0$ for any perverse sheaf $\mathcal{P}$ on $X$.
To prove this, we introduce a vanishing cycle functor of multivalued one-forms and apply techniques from non-abelian Hodge theory, both in archimedean and non-archimedean settings. These techniques allow us to deduce the desired positivity from the geometric properties of pure and mixed period maps.
@@ 1. Introduction
In the 1930s, Hopf conjectured that a closed $2n$-dimensional compact Riemannian manifold $X$ with non-positive sectional curvature satisfies $(-1)^{n}\chi(X)\geq 0$ (see [Yau82, Problem 10]). This conjecture follows from the Gauss-Bonnet theorem when $n=2$ (see [Che55]), but it remains widely open in higher dimensions. By the Cartan-Hadamard theorem, a compact Riemannian manifold $X$ with non-positive sectional curvature is aspherical, i.e., the universal cover of $X$ is contractible. Thurston, therefore, proposed the following generalization of Hopf’s conjecture (see [Kir97, Problem 4.10] and [Dav84]).
Conjecture 1.1 (Chern-Hopf-Thurston).
Let $X$ be a closed $2n$-dimensional manifold. If $X$ is aspherical, then
@@
In this general form, the conjecture is not even known in dimension 4. Charney and Davis also considered the Chern-Hopf-Thurston conjecture for piecewise Euclidean manifolds, which lead to the Charney-Davis conjecture about flag complexes (see [CD95]). Singer proposed to study the Chern-Hopf-Thurston conjecture using $L^{2}$-cohomology groups of the universal cover of $X$, leading to a stronger conjecture also involving odd-dimensional manifolds: if $X$ is a closed aspherical manifold of dimension $d$, then $b^{(2)}_{k}(\widetilde{X})=0$ for $k\neq d/2$, where $\widetilde{X}$ is the universal cover of $X$ (see [Lüc02, Chapter 11] for an overview).
Assuming $X$ is a complex projective manifold, several advances have been made (see [Gro91, JZ00, CX01, DCL24, LMW21, AW21, LIP24], among others). In [LMW21], it was observed that for projective manifolds, a natural generalization of asphericity is the property of having large fundamental group. Recall that a projective manifold $X$ is said to have large fundamental group if for any irreducible subvariety $Z\subset X$, the image of the composition $\pi_{1}(Z_{\operatorname{norm}})\to\pi_{1}(Z)\to\pi_{1}(X)$ is infinite, where $Z_{\operatorname{norm}}$ denotes the normalization of $Z$. Furthermore, the assertion that $(-1)^{n}\chi(X)\geq 0$ was strengthened to
@@
These ideas were further pursued in [AW21]. Although not explicitly stated, the following conjecture was expected to hold.
@@
In [AW21], the conjecture was proven when there exists a semisimple, almost faithful, cohomologically rigid representation $\pi_{1}(X)\to\mathrm{GL}(r,\mathbb{C})$. Here we recall that a linear representation is called almost faithful if the kernel of the representation is finite, and it is called cohomologically rigid, if it has no nontrivial first order deformation.
In this paper, we prove 1.2 within a much broader framework: we require only the existence of an almost faithful linear representation $\pi_{1}(X)\to\mathrm{GL}(r,K)$ over a field $K$ of arbitrary characteristic, without assuming semisimplity or cohomological ridigity of the representation.
@@
Let $X$ be a projective manifold. If there exists a large representation $\rho:\pi_{1}(X)\to\mathrm{GL}(r,K)$ where $K$ is any field, then $\chi(X,{\mathcal{P}})\geq 0$ for any perverse sheaves ${\mathcal{P}}$ on $X$. In particular, $(-1)^{\dim X}\chi(X)\geq 0$.
Notice that if $X$ has large fundamental group and $\rho$ is almost faithful, then $\rho$ is large. Therefore, the following corollary follows immediately.
@@
Let $X$ be a projective manifold with large fundamental group (e.g. $X$ is aspherical). If there exists an almost faithful representation $\rho:\pi_{1}(X)\to\mathrm{GL}(r,K)$, then for any perverse sheaf ${\mathcal{P}}$ on $X$, $\chi(X,{\mathcal{P}})\geq 0$. In particular, Chern-Hopf-Thurston conjecture holds for $X$.
A slightly stronger formulation of Theorem 1.3 is given below.
@@ 1.1. Sketch of the proof
We will use techniques from the linear Shafarevich conjecture presented in [DYK23, DY24, Eys04, EKPR12] to establish Theorem 1.5. As in [AW21], Arapura and the second-named author proved Theorem 1.5 assuming $\rho$ underlies a complex variation of Hodge structure ($\mathbb{C}$-VHS) with discrete monodromy. When $\rho$ does not satisfy this property, Yamanoi and the first-named author demonstrated in [DYK23] that that nontrivial multivalued one-forms can be produced on $X$. Our approach iteratively utilizes both multivalued one-forms and the period maps of $\mathbb{C}$-VHS to deduce the inequality (2).
Let us briefly explain our strategy. First, we observe that if there is a $d$-valued one-form whose image in $T^{*}X$ intersects $T^{*}_{Z}X$ at finitely many points, then the inequality (2) holds, since $d$ times the left-hand side is equal to the number of intersection points counting multiplicity.
In general, we introduce a vanishing cycle functor $\Phi_{\eta}$ of a $d$-valued one-form $\eta$ acting on the free abelian group generated by conormal varieties. We show that $d\,T^{*}_{Z}X\cdot T^{*}_{X}X=\Phi_{\eta}(T^{*}_{Z}X)\cdot T^{*}_{X}X$. Using $\Phi_{\eta}$, we can reduce inequality (2) to the same inequality with $Z$ replaced by smaller subvarieties, provided the restriction of $\eta$ to the smooth locus of $Z$ is non-trivial. The proof of Theorem 1.5 is divided into the following three cases.
Case (i): when $\mathrm{char}(K)>0$. In this case, we can prove that the above reduction process continues until $Z$ becomes a set of points. When $Z$ is a point, the inequality (2) is obvious.
Case (ii): when $K=\mathbb{C}$ and $\rho$ is semisimple. In this case, when the above reduction process terminates, $Z_{\operatorname{norm}}$ underlies a $\mathbb{C}$-VHS with large monodromy representation and discrete monodromy group. As proved in [AW21], $Z_{\operatorname{norm}}$ then admits a finite morphism to the period domain of such $\mathbb{C}$-VHS. The inequality (2) then can be deduced from the curvature properties of the period domain.
Case (iii): when $K=\mathbb{C}$ and $\rho$ is not semisimple. In this case, when the reduction process terminates, $Z_{\operatorname{norm}}$ admits a real variation of mixed Hodge structure ($\mathbb{R}$-VMHS), where its mixed period map $Z_{\rm norm}^{\rm univ}\to{\mathscr{M}}$ has discrete fibers. In this case, two additional difficulties arise compared to Case (ii):
@@
  • The mixed period domain does not have the desired non-positive curvature as in the pure case.
These difficulties are resolved by a technical result (see Proposition 5.3) which enables us to simultaneously explore the geometry of the mixed period map of the $\mathbb{R}$-VMHS and the period map of the $\mathbb{C}$-VHS corresponding to the semi-simplification of the $\mathbb{R}$-VMHS.
In [AW21], the discreteness of the monodromy group is obtained from a deep theorem of Esnault-Groechenig ([EG18], see also [Esn23] for a survey). In this paper, the discreteness of monodromy groups is deduced from nonarchemidean Hodge theory, which is same as the approach in [BKT13].
In theory, we could bypass Case (ii) and directly prove the more general Case (iii). However, we chose to present the proof in separate cases to isolate the technical arguments and to accommodate readers who are satisfied with understanding the proof up to Case (ii).
@@ 1.2. Relation with the Shafarevich conjecture and hyperbolicity
    when $X$ is not necessarily smooth and $\pi_{1}(X)$ admits a faithful reductive linear representation over $\mathbb{C}$, by the first-named author, Yamanoi and Katzarkov in [DYK23];
@@
The proofs of the Shafarevich conjecture utilize non-abelian Hodge theories both in the archimedean setting by Simpson [Sim92] and in the non-archimedean setting by Gromov-Schoen [GS92].
In this paper, we do not use any results about holomorphic convexity or Steinness. However, we do employ both archimedean and non-archimedean non-abelian Hodge theories, as used in the proofs of the Shafarevich conjectures. These techiniques are robust tools for studying the Shafarevich conjecture, the hyperbolicity of algebraic varieties (cf. e.g. [Yam10, CDY22, DY24]), and the Chern-Hopf-Thurston conjecture in the linear case. Our main result Theorem 1.5 involves a specific positivity property of the cotangent bundle of $X$, which is known to be related to the hyperbolicity of $X$. For example, a projective manifold with ample cotangent bundle is Kobayashi hyperbolic.
Although our result relies on $X$ to have large fundamental group, we hope that positivity (2) can be directly related to hyperbolicity. For example, does inequality (2) always hold for irreducible subvarieties $Z$ in a Kobayashi hyperbolic projective manifold $X$? In general, it is also interesting to find sufficient conditions on a representation of $\pi_{1}(X)$ such that inequality (2) is always strict inequality. In [CDY22, DY24], it was proved that when the Zariski closure of a (generically) large representation $\varrho:\pi_{1}(X)\to\mathrm{GL}(r,K)$ is a semisimple algebraic group, then $X$ is “almost” hyperbolic. This condition is sharp, as the case of abelian varieties must be excluded. Hence we conjecture that inequality (2) is strict if the Zariski closure of $\varrho({\rm Im}[\pi_{1}(Z_{\rm norm})\to\pi_{1}(X)])$ is a semisimple algebraic group for every positive-dimensional subvariety $Z$ of $X$. In this case, $X$ is Kobayashi hyperbolic (see [Yam10, CDY22, DY24]).
Recently, various generalizations of the Chern-Hopf-Thurston and the Singer conjectures for algebraic varieties have been formulated, e.g., in the singular setting [Max23], and in the coherent setting [AMW23]. Here, we propose Chern-Hopf-Thurston type conjectures for quasi-projective varieties. For example, we conjecture that if the universal cover of a quasi-projective manifold $X$ is a bounded symmetric space, then $\chi(X,{\mathcal{P}})\geq 0$ for any perverse sheaf on $X$ (with respect to an algebraic stratification). A special case of this conjecture is proved in the recent work [AP24].
@@ Acknowledgement
This paper builds upon previous collaborations of the second author with Yongqiang Liu, Laurentiu Maxim, and Donu Arapura, as well as the collaboration of the first author with Katsutoshi Yamanoi. We are grateful for the numerous enlightening conversations that have played a crucial role in the development of this work. We thank Jörg Schürmann for carefully reviewing an earlier draft of this paper and providing valuable feedback. We also thank Hélène Esnault for patiently answering our questions. Additionally, we have benefited greatly from discussions with Jie Liu, Qizheng Yin, and Xiping Zhang. The second author thanks Peking University and BICMR for the generous hospitality and support during the writing of this paper. The first author is partially supported by the French Agence Nationale de la Recherche (ANR) under reference ANR-21-CE40-0010 (KARMAPOLIS). The second author is partially supported by a Simons fellowship.
@@ 2. Preliminaries
We begin by reviewing some results about constructible functions, constructible sheaves, characteristic cycles and the global index formula. The recent survey article [MS22] gives a more detailed summery of these topics.
@@ 2.1. Constructible functions
is well-defined when the above stratification can be chosen to be finite and each stratum is homotopy equivalent to a finite CW-complex. For example, this holds when $X$ is compact or when $X$ is an algebraic variety and each stratum is a locally closed algebraic subvariety.
@@ 2.2. Constructible sheaves and characteristic cycles
where $K_{0}(D_{c}^{b}(X,K))$ is the Grothendieck group of $D_{c}^{b}(X,K)$. We will not go into its precise definition here (see e.g., [Dim04, Definition 4.3.19] and [KS90, Chapter IX]). Instead, we will focus on the following two important properties.
@@
Note that even though $T^{*}X$ is not compact, the intersection number is well-defined because it can be interpreted as a zero cycle in the support of ${\mathcal{F}}$ in $X$ (see [Ful98, Chapter 6]).
@@ 2.4. Multivalued one-forms
Definition 2.7 (Multivalued one-form).
Let $X$ be a complex manifold and let $E$ be a holomorphic vector bundle on $X$. A multivalued section $\eta$ on $X$ is a formal sum $\sum_{i=1}^{m}n_{i}\Gamma_{i}$, where each $n_{i}\in\mathbb{Z}_{>0}$ and each $\Gamma_{i}$ is an irreducible closed subvariety of $E$ such that the natural projection $\Gamma_{i}\to X$ is a finite surjective morphism. Such multisection $\eta$ is called $d$-valued for $d=\sum_{i=1}^{m}n_{i}\deg[\Gamma_{i}:X]$. We say that $\eta$ is reduced if all $\Gamma_{i}$ are distinct, and $\eta$ is irreducible if $m=1$ and $n_{1}=1$. We say that $\eta$ is trivial if $m=1$ and $\Gamma_{1}$ is the zero section of $E\to X$. Multivalued sections of $T^{*}X$ will be called multivalued (holomorphic) one-forms.
Let $\eta$ be a multivalued one-form on a complex manifold $X$. We sometimes denote by $\Gamma_{\eta}$ instead of $\eta$ to emphasize it is an analytic cycle in $T^{*}X$. By definition, there exists a largest Zariski open subset $X^{\circ}\subset X$ such that $\Gamma_{\eta}\cap T^{*}X^{\circ}\to X^{\circ}$ is étale. We call $X^{\circ}$ to be the unbranching locus of $\eta$ and the complement $X\setminus X^{\circ}$ the branching locus. In this case, for any $x\in X^{\circ}$, there exist an open neighborhood of $U$ of $x$ and holomorphic one-forms $\eta_{1},\ldots,\eta_{m}$ on $U$ such that $\eta|_{U}$ is equal to the union of $\eta_{1},\ldots,\eta_{m}$. Therefore, we say that $\eta$ is closed if for any point $x\in X^{\circ}$, each of the above $\eta_{i}$ is a closed one-form.
@@
Without loss of generality, we can assume that $\eta$ is irreducible. Since $X$ is projective and the natural map $\Gamma_{\eta}\to X$ is finite, $\Gamma_{\eta}$ is a projective variety. Let $Y\to\Gamma_{\eta}$ be a resolution of singularity, which is an isomorphism over the smooth locus of $\Gamma_{\eta}$. Let $\theta$ be the tautological holomorphic one-form on $T^{*}X$, and let $\theta_{Y}$ be the pullback of $\theta$ via the composition $Y\to\Gamma_{\eta}\to T^{*}X$. Since $Y$ is a projective manifold, by Hodge theory $\theta_{Y}$ must be closed. Over the unbranching locus of $\eta$, the composition $Y\to\Gamma_{\eta}\to X$ is étale, and the pushforward of $\theta_{Y}$ to $X$ is equal to $\eta$. Thus, $\eta$ is closed. ∎
@@
We will call the above map $s_{\rho}$ the (Katzarkov-Eyssidieux) reduction map for $\rho$. The construction of the multivalued holomorphic one-form $\eta_{\rho}$ associated with $\rho$ can be found in [CDY22, Step 2 in the proof of Theorem H]. The closedness of $\eta_{\rho}$ follows from Lemma 2.9. The equivalence of Item 1 and Item 3 is proved by Katzarkov [Kat97] and Eyssidieux [Eys04]. The equivalence of the first three items can be found in [CDY22, Theorem H]. Hence we only need to prove their equivalence to Item 4.